OpenStudy (anonymous):

int_{4}^{2}=-int_{1}^{4}-int_{0}^{1}+int_{0}^{2}

OpenStudy (anonymous):

well this is the original question .. Assuming that ∫[0,1,,] f(x)dx=3 ∫[0,2] f(x)dx=5 ∫[1,4] f(x)dx=8 What is ∫[4,2] f(x)dx?

OpenStudy (anonymous):

and im not sure how to go about solving it

OpenStudy (anonymous):

\[ \int_0^1f(x)dx=3, \int_0^2f(x)dx=5, \int_1^4f(x)dx=8, \int_4^2f(x)dx=? \] Is that the problem? If so: \[ \int_4^2f(x)dx=-\int_2^4f(x)dx=-\left( \int_1^4f(x)dx+\int_0^1f(x)dx-\int_0^2f(x)dx \right) \]

OpenStudy (anonymous):

So yeah, you were correct, I just misunderstood what you were saying.

OpenStudy (anonymous):

now how do i solve that part

OpenStudy (anonymous):

Just substitute in the numbers that it told you each of those integrals were equal to.

OpenStudy (anonymous):

so i plug in 3 to the first integral? a

OpenStudy (anonymous):

You plug in 3 for \(\int_0^1f(x)dx\), etc.

OpenStudy (anonymous):

thanks